### v1

@@ Abstract
An increasingly important class of estimators has members whose asymptotic distribution is non-Gaussian, yet characterizable as the $\operatorname*{arg\,max}$ of a Gaussian process. This paper presents high-level sufficient conditions under which such asymptotic distributions admit a continuous distribution function. The plausibility of the sufficient conditions is demonstrated by verifying them in three prominent examples, namely maximum score estimation, empirical risk minimization, and threshold regression estimation. In turn, the continuity result buttresses several recently proposed inference procedures whose validity seems to require a result of the kind established herein. A notable feature of the high-level assumptions is that one of them is designed to enable us to employ the celebrated Cameron-Martin theorem. In a leading special case, the assumption in question is demonstrably weak and appears to be close to minimal.
@@ 1 Introduction
An increasingly important class of estimators has members whose asymptotic distribution is non-Gaussian, yet characterizable as the $\operatorname*{arg\,max}$ of a Gaussian process. To fix ideas, letting $\bm{\theta}_{0}\in\mathbb{R}^{d}$ denote a parameter (vector) of interest, the estimators $\hat{\bm{\theta}}_{n}$ in question satisfy
@@
our goal in this paper is to give conditions on $\mu$ and $\mathcal{C}$ that imply continuity of $F_{\hat{\mathbf{s}}}$. Continuity of $F_{\hat{\mathbf{s}}}$ is useful when the goal is to use $\hat{\bm{\theta}}_{n}$ to construct confidence regions. For instance, van der Vaart (1998, Lemma 23.3) assumes continuity when establishing validity of bootstrap-based confidence intervals; see also Politis, Romano, and Wolf (1999, Section 1.2). Moreover, and relatedly, it follows from Polya’s theorem that if (1) holds and if $F_{\hat{\mathbf{s}}}$ is continuous, then the distribution functions of $r_{n}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})$ converges to $F_{\hat{\mathbf{s}}}$ not only in the bounded Lipschitz metric (or any other metric metrizing weak convergence), but also in the Kolmogorov metric; that is, we have a result of the form
$\sup_{\mathbf{t}\in\mathbb{R}^{d}}\left|\mathbb{P}\left[r_{n}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})\leq\mathbf{t}\right]-F_{\hat{\mathbf{s}}}(\mathbf{t})\right|\to 0.$ (3)
@@
In this paper we close this gap by presenting sufficient conditions for continuity of $F_{\hat{\mathbf{s}}}$ that do not require $d=1$ and are applicable (only) when $\mathcal{C}$ is not bilinear. The most notable of these conditions involves both $\mu$ and $\mathcal{C}$ and requires that for every $N\in\mathbb{N}$, restriction of $\mathcal{G}$ to $[-N,N]^{d}$ has a mean function that belongs to the reproducing kernel Hilbert space (RKHS) of its covariance kernel. By the celebrated Cameron-Martin theorem, if a Gaussian process has a mean belonging to the RKHS of its covariance kernel, then its induced probability measure and the probability measure induced by its centered version are mutually absolutely continuous. The proof of our main result uses this fact and an assumed shift equivariance property of the covariance kernel to deduce continuity of $F_{\hat{\mathbf{s}}}$.
@@
The usefulness of our main result is illustrated by applying it to three important estimation examples, namely maximum score estimation, empirical risk minimization, and threshold regression estimation. Each example involves an estimator satisfying (1) with $d$ (possibly) greater than one and a covariance kernel that is not bilinear. Although distinct in important ways, the examples enjoy the common feature that continuity of $F_{\hat{\mathbf{s}}}$ can be shown by verifying the conditions of our main result. In particular, the condition that the mean function belongs to the RKHS of the covariance kernel can be verified by following a general strategy outlined in Lifshits (1995).
@@
Although the main motivation for our paper is statistical, namely to provide easy-to-verify sufficient conditions for a continuity property needed to justify large-sample inference based on distributional approximations of the form (1), our paper also offers a contribution to the probability literature on the distributional properties of the $\operatorname*{arg\,max}$ of a Gaussian process. That problem is substantially different from the well-studied problem of understanding the distributional properties of the maximum itself, where the $d=1$ case is mostly settled (Lifshits, 1995, and references therein) and the multidimensional case is fairly well understood (Azaïs and Wschebor, 2005, and references therein). Indeed, as noted by Samorodnitsky and Shen (2013b, p. 3494), “very little is known about the random location of the supremum” of a Gaussian process.
In addition to providing results of the form (1), Kim and Pollard (1990) gave sufficient conditions for almost sure uniqueness of the $\operatorname*{arg\,max}$ of a multivariate Gaussian process; for more recent contributions along these lines, see Cox (2020) and López and Pimentel (2018) and references therein. In the case of $d=1$, moments (and sometimes continuity) of the distribution of the $\operatorname*{arg\,max}$ of a stochastic process have been studied in special cases by Groeneboom et al. (2015), Pitman and Tang (2018), Samorodnitsky and Shen (2012, 2013a, 2013b), Shen (2018), and references therein, and more recently Cattaneo et al. (2024, Lemma A.2) gave fairly general conditions guaranteeing its continuity. For $d>1$, there is scarce literature investigating the properties of the distribution of the location of the maximizer of a Gaussian process. We are only aware of Azaïs and Chassan (2020), which shows that the distribution admits a density under the assumption that the sample paths are twice differentiable. Our paper appears to be the first to provide sufficient conditions for continuity of the distribution of the $\operatorname*{arg\,max}$ of a multidimensional Gaussian process without assuming differentiability of its sample paths. Furthermore, as we discuss in Section 4, our sufficient conditions are demonstrably weak and appear to be close to minimal in a leading special case.
@@
The remainder of the paper proceeds as follows. Section 2 introduces the three above-mentioned estimation examples. Our main result is presented in Section 3, while Section 4 outlines a general strategy for verifying the conditions of the main result and demonstrates how to apply it in the examples. Finally, Section 5 compares our results with the known results alluded to in the third paragraph of this section.
@@ 2 Motivating Examples
The class of estimators satisfying (1) is rich, containing important examples in econometrics, statistics, and many other data science disciplines. To further motivate our work, this section presents three representative examples.
@@ 2.1 Maximum Score
where $\mathbbm{1}\{\cdot\}$ denotes the indicator function, $w,u\in\mathbb{R}$ and $\mathbf{x}\in\mathbb{R}^{d}$ are random variables, and $\bm{\theta}_{0}\in\Theta\subseteq\mathbb{R}^{d}$ is the parameter of interest. Manski (1975) introduced the celebrated maximum score estimator of $\bm{\theta}_{0}$, which is any maximizer $\hat{\bm{\theta}}_{n}$ of
@@
When $d=1$, because $\mu$ is quadratic and $\mathcal{C}$ is a scalar multiple of $\mathcal{C}_{\mathtt{BM}}$, it follows from van der Vaart and Wellner (1996, Exercise 3.2.5) that the distribution of $\hat{\mathbf{s}}$ is that of a scalar multiple of a random variable with a well-known continuous distribution, namely the Chernoff (1964) distribution. For $d>1$, on the other hand, it would appear to be an open question whether $F_{\hat{\mathbf{s}}}$ is continuous. We provide an affirmative answer to that question below, hereby buttressing the inference procedures for maximum score estimators proposed by Cattaneo, Jansson, and Nagasawa (2020), Delgado, Rodriguez-Poo, and Wolf (2001), Hong and Li (2020), Jun, Pinkse, and Wan (2015), Lee and Yang (2020), and Patra, Seijo, and Sen (2018) because estimated or bootstrapped quantiles of the asymptotic distribution can be shown to be consistent.
@@ 2.3 Threshold Regression
$y=\mathbf{x}^{\prime}\bm{\beta}_{0}+\mathbf{x}^{\prime}\bm{\delta}_{n}\mathbbm{1}\{q>\mathbf{w}^{\prime}\bm{\theta}_{0}\}+u,$
where $y\in\mathbb{R}$ is a dependent variable, $\mathbf{x}\in\mathbb{R}^{k}$ is a (possibly) vector-valued regressor, $q$ is a threshold variable, $\mathbf{w}\in\mathbb{R}^{d}$ is a (possibly) vector-valued factor governing the threshold cutoff, and where $\bm{\delta}_{n}$ is a “threshold effect” whose magnitude vanishes with $n$. The present model (as well as distinct generalizations thereof) has been studied by Lee, Liao, Seo, and Shin (2021) and Yu and Fan (2021), and differs from the model considered in the seminal work of Hansen (2000) by allowing the factor $\mathbf{w}$ to be non-constant.
@@
As a by-product, the continuity property of the limiting distribution function allows us conclude that if $\|\bm{\delta}_{n}\|\rightarrow 0$ and if $n\|\bm{\delta}_{n}\|^{2}\rightarrow\infty$, then
$\sup_{t\in\mathbb{R}^{d}}\left|\mathbb{P}\left[r_{n}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})\leq\mathbf{t}\right]-\mathbb{P}\left[\operatorname*{arg\,max}_{\mathbf{s}\in\mathbb{R}^{d}}\mathcal{G}(\mathbf{s};\bar{\bm{\delta}}_{n})\leq\mathbf{t}\right]\right|\to 0$
whether or not $\bar{\bm{\delta}}_{n}$ is convergent in $\mathbb{S}^{d-1}$.
@@ 3 Main Result
As before, let $\mathcal{G}$ be a Gaussian process on $\mathbb{R}^{d}$ with mean function $\mu$ and covariance kernel $\mathcal{C}$. Also, for any $N\in\mathbb{N}$, let $\mathcal{G}_{N}$ be the restriction of $\mathcal{G}$ to $[-N,N]^{d}$, let $\mu_{N}$ and $\mathcal{C}_{N}$ be the mean function and covariance kernel of $\mathcal{G}_{N}$, and let $\mathscr{H}_{N}$ be the RKHS of $\mathcal{C}_{N}$ (as defined in Section 2.6.1 of Giné and Nickl, 2016).
@@
    For any $\mathbf{s},\mathbf{t}\in\mathbb{R}^{d}$ and any $\mathbf{h}\in\mathbb{R}^{d}\backslash\{\mathbf{0}\}$, $\mathcal{C}(\mathbf{h},\mathbf{h})>0$ and
@@
Part (i) guarantees existence (with probability one) of a maximizer of $\mathcal{G}$ over $\mathbb{R}^{d}$, and over any compact set $S\subset\mathbb{R}^{d}$. By Kim and Pollard (1990, Lemma 2.6), these maximizers are unique provided $\mathbb{V}[\mathcal{G}(\mathbf{s})-\mathcal{G}(\mathbf{t})]\neq 0$ for every $\mathbf{s}\neq\mathbf{t}$. Part (ii) gives sufficient conditions for this non-degeneracy condition to hold and furthermore ensures that the centered process $\mathcal{G}^{\mu}=\mathcal{G}-\mu$ is shift equivariant in the sense that the law of the process $\mathcal{G}^{\mu}(\mathbf{h}+\cdot)-\mathcal{G}^{\mu}(\mathbf{h})$ if the same for every $\mathbf{h}\in\mathbb{R}^{d}$. By the Cameron-Martin theorem (e.g., Theorem 2.6.13 of Giné and Nickl, 2016), part (iii) ensures that the probability measures associated with $\mathcal{G}_{N}$ and $\mathcal{G}^{\mu}_{N}=\mathcal{G}_{N}-\mu_{N}$ are mutually absolutely continuous for any $N\in\mathbb{N}$, a property that plays a key role in the proof of the following result.
Theorem 1.
Under Assumption 1, $F_{\hat{\mathbf{s}}}$ in (2) is continuous.
3.1 Proof of Theorem 1
For any compact set $S\subset\mathbb{R}^{d}$, parts (i) and (ii) of Assumption 1 guarantee existence (with probability one) of a unique maximizer of $\mathcal{G}^{\mu}$ over $S$. This observation will be used repeatedly without further mention. The proof further utilizes the following implication of Assumption 1 (i)(ii).
@@ 3.1 Proof of Theorem 1
Suppose Assumption 1 (i)(ii) holds. For any $\mathbf{h}\in\mathbb{R}^{d}$, any measurable set $T\subseteq\mathbb{R}^{d}$, and any compact set $S\subset\mathbb{R}^{d}$,
@@
The (joint) distribution function of $\hat{\mathbf{s}}=(\hat{s}_{1},\dots,\hat{s}_{d})^{\prime}$ is continuous if and only if each of its marginal distribution functions is continuous. Fixing $\ell\in\{1,\dots,d\}$ and $t\in\mathbb{R}$, the proof can therefore be completed by showing that
@@
Defining $N_{t}=\lceil|t|\rceil+1$, letting $\mathbf{e}_{\ell}$ denote the $\ell$th standard basis vector of $\mathbb{R}^{d}$, and noting that
@@
$=\frac{1}{J}\mathbb{P}\left[\mathbf{e}_{\ell}^{\prime}\operatorname*{arg\,max}_{\mathbf{s}\in\underline{S}}\mathcal{G}^{\mu}(\mathbf{s})\in\left\{t+\frac{1}{2}\frac{j-1}{J-1}:1\leq j\leq J\right\}\right]$
$\leq\frac{1}{J},$
where the first equality uses Lemma 1 and the second equality uses uniqueness of the maximizer of $\mathcal{G}^{\mu}$ over $\underline{S}$. Since $J\geq 2$ was arbitrary, (7) follows. ∎
Remark.
In the proof, $N_{t}$ can be replaced by any (natural) number exceeding it and the fractions $\{(j-1)/(J-1):1\leq j\leq J\}$ appearing in the definition of the sets $\{S_{j}\}$ (and elsewhere) can be replaced by any $J$ distinct members of the unit interval.

### v2

@@ Abstract
Certain extremum estimators have asymptotic distributions that are non-Gaussian, yet characterizable as the distribution of the $\operatorname*{arg\,max}$ of a Gaussian process. This paper presents high-level sufficient conditions under which such asymptotic distributions admit a continuous distribution function. The plausibility of the sufficient conditions is demonstrated by verifying them in three examples, namely maximum score estimation, empirical risk minimization, and threshold regression estimation. In turn, the continuity result buttresses several recently proposed inference procedures whose validity seems to require a result of the kind established herein. A notable feature of the high-level assumptions is that one of them is designed to enable us to employ the Cameron-Martin theorem. In a leading special case, the assumption in question is demonstrably weak and appears to be close to minimal.
@@ 1 Introduction
Certain extremum estimators have asymptotic distributions that are non-Gaussian, yet characterizable as the distribution of the $\operatorname*{arg\,max}$ of a Gaussian process. To fix ideas, letting $\bm{\theta}_{0}\in\mathbb{R}^{d}$ denote a parameter (vector) of interest, the estimators $\hat{\bm{\theta}}_{n}$ in question satisfy
@@
our goal in this paper is to give conditions on $\mu$ and $\mathcal{C}$ that imply continuity of $F_{\hat{\mathbf{s}}}$. Continuity of $F_{\hat{\mathbf{s}}}$ is useful when the goal is to use $\hat{\bm{\theta}}_{n}$ to construct confidence regions. For instance, van der Vaart (1998, Lemma 23.3) assumes continuity when establishing validity of bootstrap-based confidence intervals; see also Politis, Romano, and Wolf (1999, Section 1.2). Moreover, and relatedly, it follows from Polya’s theorem that if (1) holds and if $F_{\hat{\mathbf{s}}}$ is continuous, then the the probability laws of $r_{n}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})$ converge to the law with distribution function $F_{\hat{\mathbf{s}}}$ not only in the bounded Lipschitz metric (or any other metric metrizing weak convergence), but also in the Kolmogorov metric; that is, we have a result of the form
$\sup_{\mathbf{t}\in\mathbb{R}^{d}}\left|\mathbb{P}\Big[r_{n}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})\leq\mathbf{t}\Big]-F_{\hat{\mathbf{s}}}(\mathbf{t})\right|\to 0.$ (3)
@@
In this paper we close this gap by presenting sufficient conditions for continuity of $F_{\hat{\mathbf{s}}}$ that do not require $d=1$ and are applicable (only) when $\mathcal{C}$ is not bilinear. Proceeding under the assumption that $d=1$, the proof of Cattaneo et al. (2024, Lemma A.2) establishes continuity of $F_{\hat{\mathbf{s}}}$ by showing that the distribution of $\hat{\mathbf{s}}$ is atomless (under the additional assumptions of the lemma). The method of proof can be adapted to give conditions under which the distribution of $\hat{\mathbf{s}}$ is atomless also when $d>1$, but when $d>1$ a distribution can be atomless even if the associated distribution function is discontinuous. Establishing continuity of $F_{\hat{\mathbf{s}}}$ when $d>1$ therefore requires a fundamentally different method of proof than that employed by Cattaneo et al. (2024, Lemma A.2). The differences in proof strategies are reflected also in the assumptions under which the proofs proceed. Notably, one of the conditions imposed in this paper explicitly involves both $\mu$ and $\mathcal{C}$ and requires that for every $N\in\mathbb{N}$, restriction of $\mathcal{G}$ to $N,N]^{d}$ has a mean function that belongs to the reproducing kernel Hilbert space (RKHS) of its covariance kernel. By the Cameron-Martin theorem, if a Gaussian process has a mean belonging to the RKHS of its covariance kernel, then its induced probability measure and the probability measure induced by its centered version are mutually absolutely continuous. The proof of our main result uses this fact and an assumed shift equivariance property of the covariance kernel to deduce continuity of $F_{\hat{\mathbf{s}}}$.
@@
The usefulness of our main result is illustrated by applying it to three examples: maximum score estimation, empirical risk minimization, and threshold regression estimation. Each example involves an estimator satisfying (1) with $d$ possibly greater than one and a covariance kernel that is not bilinear. Although distinct in several ways, the examples enjoy the common feature that continuity of $F_{\hat{\mathbf{s}}}$ can be shown by verifying the conditions of our main result. In particular, the condition that the mean function belongs to the RKHS of the covariance kernel can be verified by following a general strategy outlined in Lifshits (1995).
@@
In addition to facilitating the justification of certain large-sample inference procedures based on distributional approximations of the form (1), our paper sheds new light on the canonical problem of characterizing the distributional properties of the $\operatorname*{arg\,max}$ of a Gaussian process. That problem is substantially different from the well-studied problem of understanding the distributional properties of the maximum itself, where the $d=1$ case is mostly settled (e.g., Lifshits, 1995, and references therein), the multidimensional case is fairly well understood (e.g., Azaïs and Wschebor, 2005, and references therein), and where, more generally, continuity of the distribution function of the maximum can be established with the help of anti-concentration results (e.g., Chernozhukov, Chetverikov, and Kato, 2015). However, as noted by Samorodnitsky and Shen (2013, p. 3494), “very little is known about the random location of the supremum” of a Gaussian process. In the multidimensional case, we are only aware of Azaïs and Chassan (2020), which shows that the distribution admits a density under the assumption that the sample paths are twice differentiable.
@@
The remainder of the paper proceeds as follows. Section 2 introduces our three examples. Our main result is presented in Section 3, while Section 4 outlines a general strategy for verifying the conditions of the main result and demonstrates how to apply it in the examples. Finally, Section 5 compares our results with the known results alluded to in the third paragraph of this section.
@@ 2 Motivating Examples
The class of estimators satisfying (1) is rich, containing examples in econometrics, statistics, and other data science disciplines. To further motivate our work, this section presents three representative examples.
@@ 2.1 Maximum Score
where $\mathbbm{1}\{\cdot\}$ is the indicator function, $w,u\in\mathbb{R}$ and $\mathbf{x}\in\mathbb{R}^{d}$ are random variables, and $\bm{\theta}_{0}\in\Theta\subseteq\mathbb{R}^{d}$ is the parameter of interest. Manski (1975) introduced the maximum score estimator of $\bm{\theta}_{0}$, which is any maximizer $\hat{\bm{\theta}}_{n}$ of
@@
When $d=1$, because $\mu$ is quadratic and $\mathcal{C}$ is a scalar multiple of $\mathcal{C}_{\mathtt{BM}}$, it follows from van der Vaart and Wellner (2023, Exercise 3.2.5) that the distribution of $\hat{\mathbf{s}}$ is that of a scalar multiple of a random variable with a well-known continuous distribution, namely the Chernoff (1964) distribution. For $d>1$, on the other hand, it would appear to be an open question whether $F_{\hat{\mathbf{s}}}$ is continuous. We provide an affirmative answer to that question below, hereby buttressing a variety of inference procedures based on the maximum score estimator.
For specificity, consider the procedure of Cattaneo, Jansson, and Nagasawa (2020). That paper proposed a bootstrap-based estimator $\tilde{\bm{\theta}}_{n}^{*}$ and gave conditions under which this estimator satisfies
$r_{n}(\tilde{\bm{\theta}}_{n}^{*}-\hat{\bm{\theta}}_{n})\rightsquigarrow_{\mathbb{P}}\operatorname*{arg\,max}_{\mathbf{s}\in\mathbb{R}^{d}}\mathcal{G}(\mathbf{s}),$
where $\rightsquigarrow_{\mathbb{P}}$ denotes weak convergence in probability. Because $F_{\hat{\mathbf{s}}}$ is continuous, the displayed result can be combined with (1) to yield a bootstrap consistency result of the form
$\sup_{\mathbf{t}\in\mathbb{R}^{d}}\left|\mathbb{P}_{n}^{*}\left[\tilde{\bm{\theta}}_{n}^{*}-\hat{\bm{\theta}}_{n}\leq\mathbf{t}\right]-\mathbb{P}\left[\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\leq\mathbf{t}\right]\right|\to_{\mathbb{P}}0,$
where $\mathbb{P}_{n}^{*}$ is the bootstrap probability measure. As a consequence, for any $\bm{\lambda}\in\mathbb{R}^{d}$, defining
$q_{\bm{\lambda},n}^{*}(t)=\inf\left\{q\in\mathbb{R}:\mathbb{P}_{n}^{*}[\bm{\lambda}^{\prime}\tilde{\bm{\theta}}_{n}^{*}-\bm{\lambda}^{\prime}\hat{\bm{\theta}}_{n}\leq q]\geq t\right\},\qquad t\in(0,1),$
van der Vaart (1998, Lemma 23.3) shows that the equal-tailed “percentile” interval
$\mathsf{CI}_{\bm{\lambda},n}^{*}(1-\alpha)=\left[\bm{\lambda}^{\prime}\hat{\bm{\theta}}_{n}-q_{\bm{\lambda},n}^{*}(1-\alpha/2)~,~\bm{\lambda}^{\prime}\hat{\bm{\theta}}_{n}-q_{\bm{\lambda},n}^{*}(\alpha/2)\right]$
is a confidence interval (for $\bm{\lambda}^{\prime}\bm{\theta}_{0}$) of asymptotic level $1-\alpha$:
$\lim_{n\to\infty}\mathbb{P}\left[\bm{\lambda}^{\prime}\bm{\theta}_{0}\in\mathsf{CI}_{\bm{\lambda},n}^{*}(1-\alpha)\right]=1-\alpha.$
With minor modifications, similar comments apply to the inference procedures proposed by Delgado, Rodriguez-Poo, and Wolf (2001), Hong and Li (2020), Jun, Pinkse, and Wan (2015), Lee and Yang (2020), and Patra, Seijo, and Sen (2018).
Collectively, the inference procedures mentioned in the previous paragraph therefore constitute asymptotically valid alternatives to inference procedures based on the smoothed maximum score estimator of Horowitz (1992). (A finite-sample inference procedure for the semiparametric binary response model has recently been proposed by Rosen and Ura (2025).)
@@ 2.3 Threshold Regression
$y=\mathbf{x}^{\prime}\bm{\beta}_{0}+\mathbf{x}^{\prime}\bm{\delta}_{n}\mathbbm{1}\{q>\mathbf{w}^{\prime}\bm{\theta}_{0}\}+u,\qquad\mathbb{E}(u|\mathbf{x},q,\mathbf{w})=0,$
where $y\in\mathbb{R}$ is a dependent variable, $\mathbf{x}\in\mathbb{R}^{k}$ is a (possibly) vector-valued regressor, $q$ is a threshold variable, $\mathbf{w}\in\mathbb{R}^{d}$ is a (possibly) vector-valued factor governing the threshold cutoff, and where, borrowing ideas from the change-point literature (e.g., Bai, 1997), $\bm{\delta}_{n}$ is a “threshold effect” whose magnitude vanishes with $n$. The present model (as well as distinct generalizations thereof) has been studied by Lee, Liao, Seo, and Shin (2021) and Yu and Fan (2021), and differs from the model considered in Hansen (2000) by allowing the factor $\mathbf{w}$ to be non-constant.
@@
As a by-product, continuity of the limiting distribution function can be used to show that if $\|\bm{\delta}_{n}\|\rightarrow 0$ and if $n\|\bm{\delta}_{n}\|^{2}\rightarrow\infty$, then (whether or not $\bar{\bm{\delta}}_{n}$ is convergent in $\mathbb{S}^{d-1}$) we have
$\sup_{t\in\mathbb{R}^{d}}\left|\mathbb{P}\left[r_{n}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})\leq\mathbf{t}\right]-\mathbb{P}\left[\operatorname*{arg\,max}_{\mathbf{s}\in\mathbb{R}^{d}}\mathcal{G}(\mathbf{s};\bar{\bm{\delta}}_{n})\leq\mathbf{t}\right]\right|\to 0.$
@@ 3 Main Result
As before, let $\mathcal{G}$ be a Gaussian process on $\mathbb{R}^{d}$ with mean function $\mu$ and covariance kernel $\mathcal{C}$. Also, for any $N\in\mathbb{N}$, let $\mathcal{G}_{N}$ be the restriction of $\mathcal{G}$ to $N,N]^{d}$, let $\mu_{N}$ and $\mathcal{C}_{N}$ be the mean function and covariance kernel of $\mathcal{G}_{N}$, and let $\mathscr{H}_{N}$ be the RKHS of $\mathcal{C}_{N}$ (as defined in Section 2.6.1 of Giné and Nickl, 2016).
@@
[…]
@@
Part (i) guarantees existence (with probability one) of a maximizer of $\mathcal{G}$ over $\mathbb{R}^{d}$, and over any compact set $S\subset\mathbb{R}^{d}$. By Kim and Pollard (1990, Lemma 2.6), these maximizers are unique provided $\mathbb{V}[\mathcal{G}(\mathbf{s})-\mathcal{G}(\mathbf{t})]\neq 0$ for every $\mathbf{s}\neq\mathbf{t}$. Part (ii) gives sufficient conditions for this non-degeneracy condition to hold and furthermore implies that the centered process $\mathcal{G}^{\mu}=\mathcal{G}-\mu$ is shift equivariant in the sense that the law of the process $\mathcal{G}^{\mu}(\mathbf{h}+\cdot)-\mathcal{G}^{\mu}(\mathbf{h})$ is the same for every $\mathbf{h}\in\mathbb{R}^{d}$. An immediate implication of shift equivariance is that $\mathbb{V}[\mathcal{G}(0)]=0$. A slightly more subtle implication is recorded in the following lemma.
@@
Suppose $\mathcal{G}$ has continuous sample paths and suppose Assumption 1(ii) holds. For any $\mathbf{h}\in\mathbb{R}^{d}$, any measurable set $T\subseteq\mathbb{R}^{d}$, and any compact set $S\subset\mathbb{R}^{d}$,
@@
By the Cameron-Martin theorem (e.g., Theorem 2.6.13 of Giné and Nickl, 2016), part (iii) of Assumption 1 ensures that the probability measures associated with $\mathcal{G}_{N}$ and $\mathcal{G}^{\mu}_{N}=\mathcal{G}_{N}-\mu_{N}$ are mutually absolutely continuous for any $N\in\mathbb{N}$. Along with Lemma 1, that property plays a key role in our proof of the following result.
Theorem 1.
Under Assumption 1, $F_{\hat{\mathbf{s}}}$ in (2) is continuous.
Proof of Theorem 1. For any compact set $S\subset\mathbb{R}^{d}$, parts (i) and (ii) of Assumption 1 guarantee existence (with probability one) of a unique maximizer of $\mathcal{G}^{\mu}$ over $S$, uniqueness being a consequence of Kim and Pollard (1990, Lemma 2.6). This observation will be used repeatedly without further mention. The (joint) distribution function of $\hat{\mathbf{s}}=(\hat{s}_{1},\dots,\hat{s}_{d})^{\prime}$ is continuous if and only if each of its marginal distribution functions is continuous. Fixing $\ell\in\{1,\dots,d\}$ and $t\in\mathbb{R}$, the proof can therefore be completed by showing that
@@
[…]
@@
$=\frac{1}{J}\mathbb{P}\left[\mathbf{e}_{\ell}^{\prime}\operatorname*{arg\,max}_{\mathbf{s}\in\underline{S}}\mathcal{G}^{\mu}(\mathbf{s})\in\left\{t+\frac{1}{2}\frac{j-1}{J-1}:1\leq j\leq J\right\}\right]\leq\frac{1}{J},$
where the first equality uses Lemma 1 and the second equality uses uniqueness of the maximizer of $\mathcal{G}^{\mu}$ over $\underline{S}$. Since $J\geq 2$ was arbitrary, (7) follows. $\square$